Grade 12Mechanics

PL explain: A ball rolls of the of the top of a stairway with a horizontal velocity of magnitude of 1.8 ms-1. The steps are 0.2 m high and 0.2 m wide. Which steps will the ball hit first ( g= 10m/S2).Answer choice A) 3rd B) 4th C) 5th D ) 6th

Aishwarya
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
Since the stairs have the same width as height, find where the horizontal and vertical distances traveled are equal.
1.8m/s * t = ½at² = 4.9m/s² * t²
solutions at t = 0 s, t = 0.37 s
x = 1.8m/s * 0.37 s = 0.66 m
suggests the fourth step.
Check the edge of the third step: t = 0.6m / 1.8m/s = 0.33 s
h = ½at² = ½ * 9.8m/s² * (0.33s)² = 0.54 m
so it does miss the corner of the third step.

Regards
Arun (askIITians forum expert)